OpenStudy (anonymous):

I do not undestand !!!!! Sniff sniff find the derivative of the function. y=x^2*e^x-2e^x+2e^x

OpenStudy (anonymous):

lorda mercy

OpenStudy (anonymous):

:]

OpenStudy (anonymous):

and this JUST HW :[

OpenStudy (anonymous):

first one requires the product rule, second two do not

OpenStudy (anonymous):

(crying)........lol

OpenStudy (anonymous):

hold the phone does it really say \[y=x^2e^x-2e^x+2e^x\]

OpenStudy (anonymous):

because \(-2e^x+2e^x=0\) is why i am asking, so this is kind of dopey

OpenStudy (anonymous):

so soory its -2xe^x

OpenStudy (anonymous):

\[y=x^2e^x-2xe^x+2e^x\]makes more sense, yes

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

step by step (inch by inch) you need the product rule \[(fg)'=f'g+g'f\] for the first two terms

OpenStudy (anonymous):

first term is \[x^2e^x\] use \((fg)'=f'g+g'f\) with \(f(x)=x^2,f'(x)=2x,g(x)=e^x, g'(x)=e^x\)

OpenStudy (anonymous):

umm in the worked out solution thing 4 this book it says to take out e^x??

OpenStudy (anonymous):

you will have an \(e^x\) in each term, yes, so you can factor it out that part is algebra

OpenStudy (anonymous):

lets get the answer written first, do the algebra second

OpenStudy (anonymous):

\[y=x^2 \times e^x-2xe^x+2e^x \] use chain rule in both, it will be \[(x^2e^x + e^x \times 2x) - (2xe^x + e^x \times 2) + 2e 6x\]

OpenStudy (anonymous):

so i could do that as the first step then?

OpenStudy (anonymous):

i wouldn't, no i would do the algebra last no need to introduce another product of course you will get the same answer

OpenStudy (anonymous):

okay then

OpenStudy (anonymous):

i would just go ahead and grind it til i find it

OpenStudy (anonymous):

lol thats funny

OpenStudy (anonymous):

first term give \[2xe^x+x^2e^x\] in any order you choose (drivers ed instructor used that line when driving a stick shift)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

can we stick to orginal plz i feel i'll get confused :[

OpenStudy (anonymous):

or if you take out e^x you get \[y=e^x(x^2 -2x+2)\] the differentiation will be \[e^x(2x -2) + (x^2 - 2x - 2)e^x\]

OpenStudy (anonymous):

second one, again a product, so again the product rule \[-2(xe^x+e^x)\]

OpenStudy (anonymous):

and third term is just \(2e^x\)

OpenStudy (anonymous):

okay so ur using the second part of the function?

OpenStudy (anonymous):

don't forget that \(-2\) for the whole product

OpenStudy (anonymous):

first or second term?

OpenStudy (anonymous):

Satellite 73 ur losing me!

OpenStudy (anonymous):

ok lets go sloooooooow

OpenStudy (anonymous):

k :]

OpenStudy (anonymous):

first term is \(x^2e^x\) right? we tackle that one first

OpenStudy (anonymous):

k

OpenStudy (anonymous):

this is a product, so we need the product rule

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

the product is \(x^2\) times \(e^x\) so in english, we take the derivative of one of them, multiply it by the other (not the derivative, just the other function) and then repeat the process in the reverse order

OpenStudy (anonymous):

it'll be 2x(e^x)+E^x(x^2) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

whew now on to \(-2xe^x\)

OpenStudy (anonymous):

Im tiring u out sorry

OpenStudy (anonymous):

pull that \(-2\) right out front not at all

OpenStudy (anonymous):

so again we have a product, this time it is \(x\) times \(e^x\) so we repeat the process as before let me know what you get

OpenStudy (anonymous):

okay so let me get this straight from the product of the first term we have 2x(e^x)+e^x(x^2) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then we take care of -2xe^2 right?

OpenStudy (anonymous):

generally it looks nicer to write \(2xe^x+x^2e^x\) but no matter, multiplication is commutative

OpenStudy (anonymous):

yes, that is next

OpenStudy (anonymous):

okay sorry so we apply the product rule to -2xe^x so it'l be -2(e^x)+e^x(-2x) right?

OpenStudy (anonymous):

yes yes yes

OpenStudy (anonymous):

i mean -2(e^x)

OpenStudy (anonymous):

\[-2xe^x-2e^x\] or any combination of those

OpenStudy (anonymous):

its -2xe^x+2E^x isn't it?

OpenStudy (anonymous):

both minus by the distributive law

OpenStudy (anonymous):

u lost me :[

OpenStudy (anonymous):

how? or why?

OpenStudy (anonymous):

\[-2xe^x\] is what you have each part in the product rule will have a \(-2\) in it

OpenStudy (anonymous):

\[-2(xe^x)\] derivative is \(-2(xe^x+e^x)=-2xe^x-2e^x\) by the distributive law

OpenStudy (anonymous):

u mean taking out a -2 right? from xe^x+e^x right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

YES! okay i get u now>

OpenStudy (anonymous):

so from there do i treat each to the product rule?

OpenStudy (anonymous):

not the last term , because it is not a product of two functions, it is just \(2e^x\) so the derivative will also be \(2e^x\)

OpenStudy (anonymous):

ok so i treat the second to the last term to the product rule right?

OpenStudy (anonymous):

we have taken care of \(x^2e^x\) and got \[x^2e^x+2xe^x\] in whatever order we write it then we took care of \(-2xe^x\) and got \(-2xe^x-2e^x\)

OpenStudy (anonymous):

yes, but we did that one already

OpenStudy (anonymous):

don't do it twice!!

OpenStudy (anonymous):

ok so from here this is all?

OpenStudy (anonymous):

yes, put it all together

OpenStudy (anonymous):

AHHHHHHHHHHHHHHHHHHHHHHHH! Thanks!

OpenStudy (anonymous):

\[x^2e^x+2xe^x-2xe^x-2e^x+2e^x\] probably some algebra will clean this up a great deal and you will end up with very little

OpenStudy (anonymous):

okay thanks so much

OpenStudy (anonymous):

yw in fact, it looks like you only end up with \(x^2e^x\) hmmm it must be late

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

not reaallly its only 9:17 pm here in Cali. how late is it where ur from?

OpenStudy (anonymous):

do you know the way to san jose?

OpenStudy (anonymous):

it is 12:20 or so, i am out